A normal distribution of scores in population has a mean of µ = 100 with σ = 20. A. What is the probability of randomly selecting a score greater than X = 110 from this population? B. If a sample of n = 25 scores is randomly selected from this population, what is the probability that the sample mean will be greater than M = 110?
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 100
std deviation   =σ= 20.0000

probability of randomly selecting a score greater than X = 110 from this population:

probability = P(X>110) = P(Z>0.5)= 1-P(Z<0.5)= 1-0.6915= 0.3085

B)

sample size       =n= 25
std error=σ=σ/√n= 4.0000

probability that the sample mean will be greater than M = 110:

probability = P(X>110) = P(Z>2.5)= 1-P(Z<2.5)= 1-0.9938= 0.0062